OpenStudy (anonymous):

what is the sum of the polynomials 4k to the second power +5k and -2k to the second power +k

OpenStudy (whpalmer4):

Is that \[4k^2+5k\]\[-2k^2+k\]---------- Just do it like it appears, adding like things together. This is really what you do when you add multi-digit numbers together, the difference being that the \(k\), \(k^2\), etc. are powers of 10, and they are understood to be there rather than written out. For example: 123 456 --- 579 is really: \[1*(10)^2+2*(10)^1+3*(10)^0\]\[4*(10)^2+5*(10)^1+6*(10)^0\]--------------------------\[5*(10)^2+7*(10)^1+9*(10)^0 = 5*100+7*10+9*1 = 500+70+9=579\]

OpenStudy (whpalmer4):

Stupid OpenStudy preview gives the impression that you'll have more space than you actually get :-( Last line ends with \(500 + 70 + 9 = 579\)